# MUDAS: Mote-scale Unsupervised Domain Adaptation in Multi-label Sound Classification

### v1

@@ Abstract
I INTRODUCTION
Unsupervised Domain Adaptation (UDA) is a powerful machine learning technique that enables models to transfer knowledge from a labeled source domain to an unlabeled target domain [1, 2, 3, 4]. In real-world applications, data distribution often varies across environments—a phenomenon known as domain shift. UDA addresses this issue by adapting models to new environments without requiring additional labeled data from the target domain. It has been widely used to improve model generalization in a variety of settings, particularly in areas where labeled data is scarce or difficult to obtain.
Existing UDA algorithms, such as AdaMatch [5], are primarily designed to handle single-label classification tasks. In these tasks, each data point is assigned exactly one label, making it relatively straightforward to train models. However, real-world scenarios, such as urban sound classification, often involve multi-label classification. In these tasks, a single instance can belong to multiple classes simultaneously, such as when an audio clip contains overlapping sounds like human voices, emergency sirens, and industrial machinery. AdaMatch and similar algorithms struggle to handle the complexity of multi-label classification, as they are optimized for single-label scenarios.
@@ I INTRODUCTION
To overcome these limitations, we introduce Mote-scale Unsupervised Domain Adaptation for Sounds (MUDAS), a novel framework specifically designed to address multi-label classification in IoT environments under unsupervised domain adaptation, and use MUDAS to solve an urban sound classification problem. The framework of MUDAS is shown in Figure 1. In MUDAS, audio data from deployed IoT devices in the target domain is processed to extract feature embeddings and generate inferences. High-confidence embeddings are stored for future classifier retraining, balancing adaptation with resource constraints.
@@
To evaluate MUDAS, we use the SONYC-UST dataset [6],[7], which contains sound recordings labeled across multiple classes and recorded at various New York City locations. We simulate real-world domain shifts by splitting the dataset by location, treating one location as the source domain and another as the target. MUDAS is benchmarked against two baselines: a lower-bound model trained only on source data and tested on target data, and an upper-bound model trained and tested on target data. Results demonstrate that MUDAS outperforms these baselines, achieving good classification performance in resource-constrained, mote-scale environments.
Figure 1: The MUDAS Framework
@@
  • We explore the deployment of MUDAS on resource-constrained edge devices by optimizing its computational and storage efficiency. For computational efficiency, MUDAS performs feature embedding inference directly on edge devices and retrains only the downstream classifier, significantly reducing processing demands. To address storage limitations, we propose an algorithm that selectively store high-confidence unlabeled target data for retraining while discarding less relevant data, ensuring efficient use of limited device storage.
  • We evaluate MUDAS extensively on the SONYC-UST dataset, which includes sound recordings from diverse New York City locations. The results demonstrate that MUDAS achieves notable improvements in classification accuracy compared to existing UDA algorithms and baseline models.
II RELATED WORKS
II-A Multi-label Sound Classification
@@
II-B Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) enables models to transfer knowledge from a labeled source domain to an unlabeled target domain [1, 2, 3, 4]. A popular approach in UDA is Domain-Adversarial Training, which uses a domain discriminator to align features between source and target domains, facilitating domain-invariant feature learning and enhancing model robustness [1]. Another technique, Maximum Mean Discrepancy (MMD), reduces the domain gap by aligning the statistical means of feature distributions between source and target domains [3, 2]. Self-Ensembling, on the other hand, employs a “teacher” model to generate pseudo-labels for target data, guiding a “student” model in adapting to the target domain without requiring labeled target data [4].
AdaMatch [5] is a state-of-the-art method for unsupervised domain adaptation (UDA) and semi-supervised learning. It achieves distribution alignment between source and target domains using batch normalization statistics and employs confidence-thresholded pseudo-labeling to enhance adaptation performance. While AdaMatch has proven highly effective for multi-class image classification, applying its techniques to other domains presents unique challenges.
@@ II-C On-Device Learning
Advancements in microcontroller technology, such as the ARM Cortex-M Series [15], STM32H7 [16], ESP32-S3 [17], and NXP i.MX RT series [18], have made it possible to deploy machine learning models in resource-constrained environments. Supporting these developments, libraries like CMSIS-NN [19], LiteRT [20], and MCUNet [21] optimize model inference for embedded systems. In addition, tools like TTE [22], POET [23], and MiniLearn [24] enhance on-device training capabilities, even for highly resource-constrained mote-scale devices.
ElasticDNN [25] and RobustDA [26] address domain adaptation on edge devices; however, they primarily target systems with significantly greater computational resources than those available on mote-scale platforms. While these technological advancements are promising, the application of UDA for multi-label sound classification in highly resource-constrained, on-device environments has seen little exploration, leaving a significant gap in current research.
III SYSTEM MODEL & PROBLEM STATEMENT
@@
Each device employs a locally-stored neural network for classification, trained and fine-tuned using labeled data collected by the devices themselves. The labels are provided by human annotators. This approach [6, 27] has been used successfully to deploy cyberphysical systems where environmental monitoring is used to guide noise enforcement or remediation [28], analyze human activities, and recognize machine failure conditions.
@@
We define the original environment as the “source domain” and the new environment as the “target domain.” Our study focuses on leveraging labeled data from the source domain and unlabeled data from the target domain to learn a model that performs well on the target domain’s data.
@@
$J=\max_{\theta}\left|\text{AUPRC}^{t}(f_{\theta}(x^{s}_{[1..k]},y^{s}_{[1..k]},x^{t}_{[1..k]}))\right|$ (1)
AUPRC (Area Under the Precision-Recall Curve) is a metric used to evaluate classification performance, particularly in scenarios with class-imbalanced datasets. In this context, the AUPRC score is calculated for the test set of the target domain, denoted as $\text{AUPRC}^{t}$. The model $f_{\theta}(x^{s}_{[1..k]},y^{s}_{[1..k]},x^{t}_{[1..k]})$ is trained using source domain data points $x^{s}_{[1..k]}$ with corresponding labels $y^{s}_{[1..k]}$, and unlabeled target domain data points $x^{t}_{[1..k]})$, across all classes 1 to $k$. It is important to note that the source domain dataset $x^{s}_{[1..k]}$ and the target domain dataset $x^{t}_{[1..k]}$ may differ in size, with $n$ and $m$ data points, respectively. The goal of this paper is to maximize $J$.
@@
IV UNSUPERVISED DOMAIN ADAPTATION IN MULTI-LABEL CLASSIFICATION
Figure 2: Flow of computation in MUDAS training
In this section, we describe the elements of MUDAS; cf. its computational flow illustrated in Figure 2. Our approach use a labeled data set from the source domain and an unlabeled data set from the target domain, both consisting of audio waveforms. We utilize a pre-trained neural network to transfer the audio waveforms into feature embeddings, which are high-level representations extracted from the datasets. For feature embedding generation, we use the SONYC-L3 [14] model that is optimized for embedding extraction and inference on mote-scale devices. We train our classifier NN based on the SONYC-L3 feature embeddings (labeled from the source domain and unlabeled from the target domain).
@@ IV UNSUPERVISED DOMAIN ADAPTATION IN MULTI-LABEL CLASSIFICATION
IV-A Pre-Processing for Acoustic Data
Mel Spectrograms In sound classification, audio signals are often transformed into mel spectrograms, which are 2D time-frequency representations resembling images. This representation enables the use of advanced image-processing techniques for analyzing audio data. However, traditional Python-based signal processing libraries are unsuitable for mote-scale devices due to their resource constraints. To address this, ARM provides CMSIS (Cortex Microcontroller Software Interface Standard)[29], a suite of optimized libraries for embedded systems. Using the FFT functions in CMSIS-DSP, raw audio waveforms can be efficiently converted into mel spectrograms, enabling on-device audio preprocessing without requiring high computational resources.
Weak/Strong Augmentation To enhance generalization, we apply weak augmentations such as time-reversing (x-axis flipping) and implement strong augmentations via SpecAugment [30], which randomly masks portions of the time and frequency axes. These strategies compel the model to learn robust, invariant features, significantly improving its ability to generalize across diverse sound domains.
Feature Embeddings After augmentation, we extract feature embeddings for urban sound classification, a method well-suited for the limited datasets in this domain. Unlike AdaMatch, which uses ResNet[31] for image-based tasks, our approach leverages embeddings specifically tailored for sound data. This not only improves generalization but also provides more interpretable representations of data relationships. By utilizing pre-trained models, we reduce the need for large datasets while maintaining efficiency through dimensionality reduction. For embedding generation, we use the SONYC-L3 model, optimized for efficient operation on resource-constrained devices.
IV-B Training Classifier Model with Feature Embeddings
@@
$z^{ss}=\lambda\cdot z^{ss^{\prime}}+(1-\lambda)\cdot z^{ss^{\prime\prime}}$ (2)
$z^{ws}=\lambda\cdot z^{ws^{\prime}}+(1-\lambda)\cdot z^{ws^{\prime\prime}}$ (3)
@@
$\tilde{y}^{wt}_{(i,j)}=min(z^{wt}_{(i,j)}\cdot\frac{\frac{1}{n}\sum_{j=1}^{n}z^{ws}_{(i,j)}}{\frac{1}{m}\sum_{j=1}^{m}z^{wt}_{(i,j)}},1)$ (4)
@@
Two source losses, $l^{ws}$ and $l^{ss}$, handle weak and strong augmentations, respectively. Complementing these, the two target $l^{+}$ for positive classes and $l^{-}$ for negative classes—refine the model’s performance on the target domain by reinforcing correct classifications and reducing uncertainty in label assignment.
$l^{ws}=-\frac{1}{k}\frac{1}{n}\sum_{i=1}^{k}\sum_{j=1}^{n}[y^{s}_{(i,j)}\cdot logz^{ws}_{(i,j)}-(1-y^{s}_{(i,j)})\cdot log(1-z^{ws}_{(i,j)})]$ (7)
$l^{ss}=-\frac{1}{k}\frac{1}{n}\sum_{i=1}^{k}\sum_{j=1}^{n}[y^{s}_{(i,j)}\cdot logz^{ss}_{(i,j)}-(1-y^{s}_{(i,j)})\cdot log(1-z^{ss}_{(i,j)})]$ (8)
$l^{+}=-\frac{1}{k}\frac{1}{m}\sum_{i=1}^{k}\sum_{j=1}^{m}[\mathds{1}(\tilde{y}^{wt}_{(i,j)}\geq c^{+}_{i})\cdot logz^{st}_{(i,j)}]$ (9)
$l^{-}=-\frac{1}{k}\frac{1}{m}\sum_{i=1}^{k}\sum_{j=1}^{m}[\mathds{1}(\tilde{y}^{wt}_{(i,j)}\leq c^{-}_{i})\cdot(1-logz^{st}_{(i,j)})]$ (10)
Diversity Regularization To further address the challenge of dataset imbalance, we apply the Diversity Regularization approach proposed in [32]. This method introduces two distinct regularization strategies—one for positive labels and another for negative labels—ensuring the model gives balanced attention to both present and absent labels. By preventing overconfidence in individual labels, these strategies encourage diverse and well-calibrated predictions, which is especially critical in multi-label classification tasks where multiple correct labels can coexist.
$l^{+}_{div}=\frac{1}{k}\frac{1}{m}\sum_{i=1}^{k}\sum_{j=1}^{m}[\mathds{1}(\tilde{y}^{wt}_{(i,j)}\geq c^{+}_{i})\cdot z^{st}_{(i,j)}logz^{st}_{(i,j)}]$ (11)
$l^{-}_{div}=\frac{1}{k}\frac{1}{m}\sum_{i=1}^{k}\sum_{j=1}^{m}[\mathds{1}(\tilde{y}^{wt}_{(i,j)}\leq c^{-}_{i})\cdot(1-z^{st}_{(i,j)})\cdot log(1-z^{st}_{(i,j)})]$ (12)
@@
V ON-DEVICE RESOURCE-EFFICIENT RELEARNING
@@
V-A Optimization in Relearning Algorithm
@@
Efficient Feature Representation Reuse We enhance MUDAS’s compute efficiency by reusing feature representations for the urban sounds context that are known to be robust and to generalize well across domains [28]. For embedding generation, we utilize the SONYC-L3 model, optimized for resource-constrained devices. This model reduces the melspectrogram’s granularity and decreases the number of convolutional filters, retaining performance close to its predecessor L3 model (whose static and dynamic memory needs exceed that of mote-scale IoT devices) and outperforming VGGish. Moreover, instead of retraining the entire model, we decompose it, focusing on generating stable feature embeddings, and only fine-tuning the downstream classifier for the target domain for adapting the model (cf. Figure 1).
@@
V-B Selection for Data Used in Relearning
Figure 3: Micro AUPRC performance of varying unlabeled target dataset sizes vs. consistent dataset size with different splits (3-class evaluation)
TABLE I: The Micro AUPRC performance of MUDAS for target data samples categorized by probability ranges: high, mid, and low (3-class evaluation)

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: The 12th ACM International Conference on Systems for Energy-Efficient Buildings, Cities, and Transportation; November 19–21, 2025; Golden, CO, USAbooktitle: The 12th ACM International Conference on Systems for Energy-Efficient Buildings, Cities, and Transportation (BuildSys ’25), November 19–21, 2025, Golden, CO, USAdoi: 10.1145/3736425.3770098isbn: 979-8-4007-1945-5/25/11ccs: Computing methodologies Online learning settingsccs: Applied computing Sound and music computing
1. INTRODUCTION
UDA is a powerful machine learning technique that enables models to transfer knowledge from a labeled source domain to an unlabeled target domain (Domain_Adversarial_1; Domain_Adversarial_2; MMD_1; Ensembling). In real-world applications, data distribution often varies across environments—a phenomenon known as domain shift. UDA addresses this issue by adapting models to new environments without requiring additional labeled data from the target domain. It has been widely used to improve model generalization in a variety of settings, particularly in areas where labeled data is scarce or difficult to obtain.
Existing UDA algorithms, such as AdaMatch (AdaMatch), are primarily designed to handle single-label classification tasks. In these tasks, each data point is assigned exactly one label, making it relatively straightforward to train models. However, real-world scenarios, such as urban sound classification, often involve multi-label classification. In these tasks, a single instance can belong to multiple classes simultaneously, such as when an audio clip contains overlapping sounds like human voices, emergency sirens, and industrial machinery.
@@ 1. INTRODUCTION
To overcome these limitations, we introduce MUDAS, a novel framework specifically designed to address multi-label classification in IoT environments under unsupervised domain adaptation, and use MUDAS to solve an urban sound classification problem. The framework of MUDAS is shown in Figure 1. In MUDAS, audio data from deployed IoT devices in the target domain is processed to extract feature embeddings and generate inferences. High-confidence embeddings are stored for future classifier retraining, balancing adaptation with resource constraints.
@@
To evaluate MUDAS, we use the SONYC-UST dataset (SONYC_UST_V1),(SONYC_UST_V2), which contains sound recordings labeled across multiple classes and recorded at various New York City locations. We simulate real-world domain shifts by splitting the dataset by location, treating one location as the source domain and another as the target. MUDAS is benchmarked against two baselines: a lower-bound model trained only on source data and tested on target data, and an upper-bound model trained and tested on target data. Results demonstrate that MUDAS outperforms these baselines, achieving good classification performance in resource-constrained, mote-scale environments.
Figure 1. The MUDAS Framework
@@
  • We explore the deployment of MUDAS on resource-constrain-ed edge devices by optimizing its computational and storage efficiency. For computational efficiency, MUDAS performs feature embedding inference directly on edge devices and retrains only the downstream classifier, significantly reducing processing demands. To address storage limitations, we propose an algorithm that selectively store high-confidence unlabeled target data for retraining while discarding less relevant data, ensuring efficient use of limited device storage.
  • We evaluate MUDAS extensively on the SONYC-UST dataset, which includes sound recordings from diverse New York City locations. The results demonstrate that MUDAS achieves notable (up to 6.7%) improvements in classification accuracy relative to the source model.
2. RELATED WORKS
2.1. Multi-label Sound Classification
@@
2.2. Unsupervised Domain Adaptation
UDA enables models to transfer knowledge from a labeled source domain to an unlabeled target domain. A popular approach in UDA is Domain-Adversarial Training, which uses a domain discriminator to align features between source and target domains, facilitating domain-invariant feature learning and enhancing model robustness (Domain_Adversarial_1). Another technique, Maximum Mean Discrepancy (MMD), reduces the domain gap by aligning the statistical means of feature distributions between source and target domains (MMD_1; Domain_Adversarial_2). Self-Ensembling, on the other hand, employs a “teacher” model to generate pseudo-labels for target data, guiding a “student” model in adapting to the target domain without requiring labeled target data (Ensembling).
AdaMatch (AdaMatch) is a state-of-the-art method for UDA and semi-supervised learning. It achieves distribution alignment between source and target domains using batch normalization statistics and employs confidence-thresholded pseudo-labeling to enhance adaptation performance. While AdaMat-ch has proven highly effective for multi-class image classification, applying its techniques to other domains presents unique challenges.
@@ 2.3. On-Device Learning
Advancements in microcontroller technology, such as the ARM Cortex-M Series (M7), STM32H7 (STM32H7), and NXP i.MX RT series (iMX_RT), have made it possible to deploy machine learning models in resource-constrained environments. Supporting these developments, libraries like CMSIS-NN (CMSIS_NN), LiteRT (LiteRT), and MCUNet (MCUNet) optimize model inference for embedded systems. In addition, tools like TTE (256KB), POET (POET), and MiniLearn (MiniLearn) enhance on-device training capabilities, even for highly resource-constrained mote-scale devices. While these technological advancements are promising, the application of UDA for multi-label sound classification in highly resource-constrained, on-device environments has seen little exploration, leaving a significant gap in current research.
3. SYSTEM MODEL & PROBLEM STATEMENT
@@
Each device employs a locally-stored neural network for classification, trained and fine-tuned using labeled data collected by the devices themselves. The labels are provided by human annotators. This approach (SONYC_UST_V1; alias2019review) has been used successfully to deploy cyberphysical systems where environmental monitoring is used to guide noise enforcement or remediation (bello2019sonyc), analyze human activities, and recognize machine failure conditions.
@@
We define the original environment as the “source domain” and the new environment as the “target domain”. Our study focuses on leveraging labeled data from the source domain and unlabeled data from the target domain to learn a model that performs well on the target domain’s data.
@@
(1) $J=\max_{\theta}\left|\text{AUPRC}^{t}(f_{\theta}(x^{s}_{[1..k]},y^{s}_{[1..k]},x^{t}_{[1..k]}))\right|$
AUPRC (Area Under the Precision-Recall Curve) is a metric used to evaluate classification performance, particularly in scenarios with class-imbalanced datasets. In this context, the AUPRC score is calculated for the test set of the target domain, denoted as $\text{AUPRC}^{t}$. The model $f_{\theta}(x^{s}_{[1..k]},y^{s}_{[1..k]},x^{t}_{[1..k]})$ is trained using source domain data points $x^{s}_{[1..k]}$ with corresponding labels $y^{s}_{[1..k]}$, and unlabeled target domain data points $x^{t}_{[1..k]}$, across all classes 1 to $k$. It is important to note that the source domain dataset $x^{s}_{[1..k]}$ and the target domain dataset $x^{t}_{[1..k]}$ may differ in size, with $n$ and $m$ data points, respectively. The goal of this paper is to maximize $J$.
@@
4. UNSUPERVISED DOMAIN ADAPTATION IN MULTI-LABEL CLASSIFICATION
Figure 2. Flow of computation in MUDAS training
In this section, we describe the elements of MUDAS; cf. its computational flow illustrated in Figure 2. Our approach use a labeled data set from the source domain and an unlabeled data set from the target domain, both consisting of audio waveforms. We utilize a pre-trained neural network to transfer the audio waveforms into feature embeddings, which are high-level representations extracted from the datasets. For feature embedding generation, we use the SONYC-L3 (MKII) model that is optimized for embedding extraction and inference on mote-scale devices. We train our classifier NN based on the SONYC-L3 feature embeddings (labeled from the source domain and unlabeled from the target domain).
@@ 4. UNSUPERVISED DOMAIN ADAPTATION IN MULTI-LABEL CLASSIFICATION
4.1. Pre-Processing for Acoustic Data
Mel Spectrograms In sound classification, audio signals are often transformed into mel spectrograms, which are 2D time-frequency representations resembling images. This representation enables the use of advanced image-processing techniques for analyzing audio data. However, traditional Python-based signal processing libraries are unsuitable for mote-scale devices due to their resource constraints. To address this, ARM provides CMSIS (Cortex Microcontroller Software Interface Standard)(CMSIS), a suite of optimized libraries for embedded systems. Using the FFT functions in CMSIS-DSP, raw audio waveforms can be efficiently converted into mel spectrograms, enabling on-device audio preprocessing without requiring high computational resources.
Weak/Strong Augmentation To enhance generalization, we apply weak augmentations such as time-reversing (x-axis flipping) and implement strong augmentations via SpecAugment (SpecAugment), which randomly masks portions of the time and frequency axes. These strategies compel the model to learn robust, invariant features, significantly improving its ability to generalize across diverse sound domains.
Feature Embeddings After augmentation, we extract feature embeddings for urban sound classification, a method well-suited for the limited datasets in this domain. Unlike AdaMatch, which uses ResNet(ResNet) for image-based tasks, our approach leverages embeddings specifically tailored for sound data. This not only improves generalization but also provides more interpretable representations of data relationships. By utilizing pre-trained models, we reduce the need for large datasets while maintaining efficiency through dimensionality reduction. For embedding generation, we use the SONYC-L3 model, optimized for efficient operation on resource-constrained devices.
4.2. Training Classifier Model with Feature Embeddings
@@
(2) $z^{ss}=\lambda\cdot z^{ss^{\prime}}+(1-\lambda)\cdot z^{ss^{\prime\prime}}$
(3) $z^{ws}=\lambda\cdot z^{ws^{\prime}}+(1-\lambda)\cdot z^{ws^{\prime\prime}}$
@@
(4) $\tilde{y}^{wt}_{(i,j)}=min(z^{wt}_{(i,j)}\cdot\frac{\frac{1}{n}\sum_{j=1}^{n}z^{ws}_{(i,j)}}{\frac{1}{m}\sum_{j=1}^{m}z^{wt}_{(i,j)}},1)$
@@
[…]
(9) $l^{+}=-\frac{1}{k}\frac{1}{m}\sum_{i=1}^{k}\sum_{j=1}^{m}[\mathds{1}(\tilde{y}^{wt}_{(i,j)}\geq c^{+}_{i})\cdot logz^{st}_{(i,j)}]$
[…]
Diversity Regularization To further address the challenge of dataset imbalance, we apply the Diversity Regularization approach proposed in (Diversity_Regularization). This method introduces two distinct regularization strategies—one for positive labels and another for negative labels—ensuring the model gives balanced attention to both present and absent labels. By preventing overconfidence in individual labels, these strategies encourage diverse and well-calibrated predictions, which is especially critical in multi-label classification tasks where multiple correct labels can coexist.
(11) $l^{+}_{div}=\frac{1}{k}\frac{1}{m}\sum_{i=1}^{k}\sum_{j=1}^{m}[\mathds{1}(\tilde{y}^{wt}_{(i,j)}\geq c^{+}_{i})\cdot z^{st}_{(i,j)}logz^{st}_{(i,j)}]$
[…]
@@
5. ON-DEVICE RESOURCE-EFFICIENT RELEARNING
@@
5.1. Optimization in Relearning Algorithm
@@
Efficient Feature Representation Reuse We enhance MUDAS’s compute efficiency by reusing feature representations for the urban sounds context that are known to be robust and to generalize well across domains (bello2019sonyc). For embedding generation, we utilize the SONYC-L3 model, optimized for resource-constrained devices. This model reduces the melspectrogram’s granularity and decreases the number of convolutional filters, retaining performance close to its predecessor L3 model (whose static and dynamic memory needs exceed that of mote-scale IoT devices) and outperforming VGGish. Moreover, instead of retraining the entire model, we decompose it, focusing on generating stable feature embeddings, and only fine-tuning the downstream classifier for the target domain for adapting the model (cf. Figure 1).
@@
Figure 3. Micro AUPRC performance of varying unlabeled target dataset sizes vs. consistent dataset size with different splits (3-class evaluation)
Table 1. The Micro AUPRC performance of MUDAS for target data samples categorized by probability ranges: high, mid, and low (3-class evaluation)